OpenStudy (anonymous):

(x^2 + 1)(x^3 + 2x)(x^2 + 4x - 16i - 4xi) = 0 Solve the following equation by identifying all of its roots including any imaginary numbers. In your final answer, include the necessary steps and calculations.

OpenStudy (anonymous):

set each factor equal to zero and solve

OpenStudy (anonymous):

for example \[x^2 + 1=0\\ x^2=-1\\ x=\pm\sqrt{-1}\\x=\pm i\] for the first factor

OpenStudy (anonymous):

\[x^2 + 4x - 16i - 4xi=0\] will present a challenge but if you rewrite it as \[x^2+(4-4i)x-16i=0\] you can use the quadratic formula with \[a=1,b=4-4i,c=-16i\]

OpenStudy (anonymous):

Thank you very much. Do I have to do anything else or is everything you wrote is the answer?

OpenStudy (anonymous):

@satellite73